OpenStudy (anonymous):

Calculus help pleasee Find any critical numbers of the function. f(theta) = 2 sec theta + tan theta , 0< theta < 2 pi

OpenStudy (anonymous):

Take the derivative of your function and set it equal to zero.

OpenStudy (anonymous):

well, I can do that.

OpenStudy (anonymous):

I'm asking for help on solving for a derivative.

OpenStudy (anonymous):

\[f(\theta) = 2\sec\theta + \tan\theta\]\[f'(\theta) = 2\sec\theta\tan\theta + \sec^2\theta \]

OpenStudy (dan815):

crit points are where the slope of the graph = 0?

OpenStudy (anonymous):

@SACAPUNTAS Yay! I got the right derivative. then you have to set that equal to 0 and solve for theta right? sorry I had to mention that earlier :(

OpenStudy (anonymous):

Yes like @dan815 said, it's where the slope (i.e. the derivative) = 0

OpenStudy (anonymous):

then do you think this step is right? after you found the derivative, I did sec theta(2tantheta + sec theta) = 0

OpenStudy (anonymous):

and I divided sec theta on both sides so it became 0 = 2 tan theta + sec theta ?

OpenStudy (anonymous):

and I was kinda stuck there.. :(

OpenStudy (jhannybean):

\[\large f'(\theta) = 2\sec(\theta)\tan(\theta) + \sec^2(\theta)=0\]\[\large f'(\theta) = \sec(\theta)(2\tan(\theta) +\sec(\theta))=0\]